OpenStudy (anonymous):

how to intergrate x/(5-x)^1/2 by using substitution method?

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ x }{ \sqrt{5-x} }dx\] what kind of a substitution are required to make ? algebraic ? trigonometric ?

OpenStudy (anonymous):

you could use x=x-5+5 and distribute the denominator between (x-5) and +5

OpenStudy (anonymous):

are't we supposed to let U= something? but i dont know which term to let

OpenStudy (anonymous):

but why 5-x becomes the numerator after you let 5-x=u?

OpenStudy (calculator):

because u let u=5-x u=5-x x=5-u same thing

OpenStudy (anonymous):

oh okay, I get it ! but my lecturer taught different way. Thanks :)